OpenStudy (anonymous):

find LCM of x^2+5x+6 and (x+3)(x-1)

OpenStudy (anonymous):

\[x ^{2}+5x+6=(x+2)(x+3)\] and second expression is (x+3)(x-1) so lcm=(x+3)(x-1)(x+2) also GCM=(x+3)

OpenStudy (anonymous):

yay! thnx very very much!